### v1

# Bine Trees: Enhancing Collective Operations by Optimizing Communication Locality

Abstract.
Communication locality plays a key role in the performance of collective operations on large HPC systems, especially on oversubscribed networks where groups of nodes are fully connected internally but sparsely linked through global connections. We present Bine (binomial negabinary) trees, a family of collective algorithms that improve communication locality. Bine trees maintain the generality of binomial trees and butterflies while cutting global-link traffic by up to $33\%$. We implement eight Bine-based collectives and evaluate them on four large-scale supercomputers with Dragonfly, Dragonfly+, oversubscribed fat-tree, and torus topologies, achieving up to $5\times$ speedups and consistent reductions in global-link traffic across different vector sizes and node counts.
collective communication, topology, binomial tree, MPI, NCCL
copyright: acmlicensedjournalyear: 2025copyright: ccconference: The International Conference for High Performance Computing, Networking, Storage and Analysis; November 16–21, 2025; St Louis, MO, USAbooktitle: The International Conference for High Performance Computing, Networking, Storage and Analysis (SC ’25), November 16–21, 2025, St Louis, MO, USAdoi: 10.1145/3712285.3759835isbn: 979-8-4007-1466-5/2025/11ccs: Computing methodologies Distributed algorithmsccs: Networks Data center networks
1. Introduction
As high-performance computing (HPC) systems and data centers continue to grow (500, 2025; Atchley et al., 2023a; Meta, [n. d.]; Morgan, 2025), communication locality plays a crucial role in determining performance. In most HPC networks, communication with nearby nodes is generally faster than with distant nodes. This is particularly evident in scale-up networks, where bandwidth can be up to an order of magnitude higher than that of the scale-out network (De Sensi et al., 2024b; Hoefler et al., 2022). However, even when focusing solely on the scale-out network, locality remains important. While full-bandwidth scale-out networks such as fat trees can deliver high performance, they also come with significant cost and power consumption. As a result, oversubscribed (tapered) networks are often used. These are typically organized into fully connected groups, such as groups in Dragonfly/Dragonfly+ topologies (Kim et al., 2008; Shpiner et al., 2017; De Sensi et al., 2020; Atchley et al., 2023b) or subtrees in oversubscribed fat trees, which are then interconnected more sparsely through global links.
Reducing the traffic forwarded over global links is advantageous for several reasons. First, global links are often longer than local links, resulting in higher communication latency (De Sensi et al., 2020; Atchley et al., 2023b). Second, because global links are typically oversubscribed, they can sustain only a limited number of concurrent communications, often causing congestion when multiple ranks from the same or different applications contend for the same link, leading to higher latency and lower throughput (Chunduri et al., 2019; De Sensi et al., 2019; Smith et al., 2018; De Sensi et al., 2024b). This issue is amplified in oversubscribed low-diameter networks, where jobs are usually distributed across multiple groups. Such distribution helps arbitrary traffic patterns approximate uniform behavior, under which these topologies achieve high throughput (Wang et al., 2025; Yang et al., 2016; Kang et al., 2024; Blach et al., 2024; SchedMD, 2025b), but it also increases global link utilization. Third, global links consume more energy than local links, making it crucial to reduce their usage (Abts et al., 2010). Last, as the physical size of data centers and supercomputers approaches practical limits, tightly coupled applications run across multiple geographically distributed data centers (Luo et al., 2024; Gemini Team, 2023; OpenAI, 2025; Dylan Patel, Daniel Nishball and Jeremie Eliahou Ontiveros, [n. d.]). In such scenarios, communication between distinct data centers or supercomputers must traverse even slower Wide-Area Network (WAN) links, and improving communication locality is critical.
For these reasons, running collectives on oversubscribed networks requires careful scheduling of communications to minimize traffic on global links and ensure scalability. Consider, for example, a broadcast operation implemented with a binomial tree. The root first sends the data to one process, and then each process forwards it to another, doubling the number of participants at each step until all processes have received the data. The algorithm remains correct as long as each process that has received the data sends it to one that has not, but different choices of destination processes can result in different amounts of traffic on global links.
For instance, consider the scenario in Fig. 1, where eight nodes are connected through a 2:1 oversubscribed fat-tree. Each leaf switch has two downlinks and one uplink, serving two nodes per switch. In a distance-doubling binomial tree (as in Open MPI (MPI, 2025a, b), top part of Fig. 1), the distance between communicating ranks doubles at each step (e.g., rank 0 first sends its data to rank 1, then to rank 2). In contrast, in a distance-halving tree (as in MPICH (MPICH, 2025a), bottom part of the Fig. 1), the distance halves at each step (e.g., rank 0 first sends its data to rank 4, then to rank 2). In both cases, each rank that has already received the $n$ bytes of the vector forwards them to another rank. The distance-doubling broadcast forwards a total of $6n$ bytes over global links ($2n$ at step $1$ and $4n$ at step $2$), compared to only $3n$ bytes for the distance-halving variant.
Figure 1. Traffic on global links for a broadcast collective using distance-doubling and distance-halving binomial trees.
[…]
Traffic on global links for a broadcast collective using distance-doubling and distance-halving binomial trees.
This shows that, despite both algorithms requiring the same number of steps and transferring the same total data volume, the distance-halving variant is more efficient in reducing global link usage. Intuitively, closer communication is preferable when more ranks are involved in sending data. Furthermore, if all links offer the same bandwidth, having two data transfers sharing a link can halve the effective injection bandwidth, as in step 1 of distance-doubling, between ranks 0-2 and 1-3. Though simplified for illustrative purposes, this example shows how the communication schedule affects inter-group traffic and performance.
Moreover, although ring algorithms reduce global link traffic compared to binomial trees and butterflies, their performance is suboptimal for large numbers of ranks and/or small to medium vector sizes (De Sensi et al., 2024a). Also, it is important to note that existing hierarchical and locality-aware collective algorithms may not always be effective. These algorithms typically require prior knowledge of the number of processes per group and often assume that jobs run with the same number of processes in each group (Bienz et al., 2022; Träff, 2006; Karonis et al., 2000; Kandalla et al., 2009; Graham et al., 2011). While such assumptions are reasonable for the number of processes per node, they may not hold when considering the partitioning of processes across subtrees or Dragonfly groups. In fact, the scheduler’s process-to-node allocation is not known in advance and is unlikely to result in an even distribution of processes across groups.
Based on these insights, this paper contributes as follows:
  • We introduce Bine (binomial negabinary) trees [^1] A bine is a twisting vine, reminding the way in which we build Bine trees (Sec. 2.2). , a novel way of constructing binomial trees (and butterflies) collective algorithms, reducing the distance between communicating processes and, consequently, the volume of data transmitted across groups (Sec. 2-3). Bine trees are topology-agnostic and generic as binomial trees and butterflies, but additionally optimize communication locality. They are orthogonal to hierarchical algorithms and can be used in both hierarchical and flat algorithms.
  • We design and implement new collective algorithms based on Bine trees and butterflies for allgather, allreduce, reduce-scatter, alltoall, broadcast, gather, reduce, scatter (Sec. 4) [^2] Reference implementations are available at https://github.com/HLC-Lab/bine-trees..
  • We evaluate Bine collectives on up to $8192$ nodes (Sec. 5), on four supercomputers with diverse network topologies: LUMI (Dragonfly), Leonardo (Dragonfly+), MareNostrum 5 (2:1 oversubscribed fat-tree), and Fugaku (torus). We include three different MPI implementations: Fujitsu MPI, Open MPI, and Cray MPICH, as well as NCCL. Bine trees consistently outperform state-of-the-art algorithms, improving performance by up to 80% on LUMI and MareNostrum 5, 50% on Leonardo, and up to $5\times$ on Fugaku, while also reducing global link load by up to 33%. These results demonstrate the effectiveness of Bine trees in optimizing collective communication, delivering substantial performance and scalability benefits across diverse networks and software stacks.
2. Distance-Halving Bine Trees
We now recall the construction of binomial trees (Sec. 2.1), then introduce distance-halving Bine trees visually (Sec. 2.2) and formally (Sec. 2.3), and discuss their advantages over binomial trees (Sec. 2.4).
2.1. Binomial Trees Construction
A distance-halving binomial tree of order 0 consists of a single node. A tree of order $k$ is built recursively by linking two trees of order k1k-1, making one the child of the other. This structure is shown in the top part of Fig. 2, where each subtree is itself a binomial tree. When used in collective operations, each node maps to a rank, and each edge represents communication between ranks. Any mapping of ranks to nodes yields a valid algorithm, also for collectives involving computation (e.g., reduce), as long as the operator is associative, which MPI assumes by default.
Figure 2. Distance-halving binomial tree construction.
[…]
Distance-halving binomial tree construction.
Although the view of binomial trees shown in the top part of Fig. 2 highlights the recursive construction of the tree, it does not clearly illustrate how communication progresses step by step. To make this explicit, we use the representation in the bottom part of Fig. 2, where ranks are arranged left to right and each row (indicated with a dashed horizontal line) corresponds to a communication step. Boxes indicate that the value is present in a rank’s memory. Vertical lines show a rank lifetime, while diagonal arrows represent message transfers between ranks. For example, in a broadcast across four ranks (i.e., a tree of order 2) with rank 0 as root, rank 0 first sends the data to rank 2 (Fig. 2, [Uncaptioned image]). In the next step, rank 0 sends the data to rank 1 (Fig. 2, [Uncaptioned image]), while rank 2 sends it to rank 3 (Fig. 2, [Uncaptioned image]).
2.2. Bine Trees Construction: Visual Intuition
The key idea behind Bine trees is to build a binomial tree by arranging ranks so that the distance between communicating ranks is smaller than in standard binomial trees. Shorter distances increase the likelihood that communication occurs within the same group (or within the same node), thereby benefiting from higher bandwidth and lower latency. Since our goal is to design an algorithm as simple and generic as standard binomial tree algorithms, and not tied to any specific topology, we do not consider the actual physical distance between nodes (e.g., by retrieving their location (Ma et al., 2011)).
In standard binomial trees, the distance between communicating ranks is defined by the difference in their identifiers. For instance, in distance-halving binomial trees, rank $6$ first communicates with rank $4$ (distance $2$) and then with rank $7$ (distance $1$). This provides a reasonable approximation of physical distance when ranks are mapped linearly across the system, which is often the default choice. For example, this is the case in Slurm’s block distribution (SchedMD, 2025a).
We make the same assumption that ranks are mapped linearly. If this is not the case, a remapping to a block distribution can be applied when the communicator is created, and the remapped rank identifiers can be used instead of the original ones. For instance, on the systems used in this paper, hostnames are numbered consecutively across groups or subtrees, so it is enough to sort ranks by hostname. However, differently from binomial trees, we consider the modulo distance between rank identifiers. Namely, we arrange the $p$ ranks in a circle $0,1,\ldots,p-1$, and consider the distance between two ranks $r$ and $q$ as the minimum distance along the circle. I.e., $d(r,q)=\min((r-q)\bmod p,\,(q-r)\bmod p)$.
We construct a distance-halving Bine tree recursively. A two-node Bine tree (Fig. 3, order 1) is identical to a standard binomial tree. To build a four-node (i.e., order 2) Bine tree, we connect an order 1 Bine tree (Fig. 3, [Uncaptioned image]), to a mirrored order 1 Bine tree (i.e., reflected on the vertical axis) and placed to its left (Fig. 3, [Uncaptioned image]). The mirrored tree is placed to the left rather than to the right to keep a shorter distance between the roots of the subtrees (one rather than two as in a binomial tree – see Fig. 2, [Uncaptioned image]). Ranks are numbered left to right starting from the root, wrapping around once the rightmost node of the tree is reached (e.g., on the 4-node tree, ranks $2$ and $3$ are those on the left of the root).
Figure 3. Distance-halving Bine trees construction.
[…]
Distance-halving Bine trees construction.
The process generalizes recursively: an eight-node Bine tree is constructed by combining an order-2 Bine tree (Fig. 3, [Uncaptioned image]) and a mirrored order-2 Bine tree (Fig. 3, [Uncaptioned image]), placed side by side and arranged so that the modulo distance between the roots is as small as possible. In this case, placing the new four-node subtree on the right (rather than on the left) yields a root-to-root modulo distance of three, compared to five in the alternative placement. For comparison, the corresponding binomial tree would result in a distance of four (Fig. 2, [Uncaptioned image]). If $p$ is not a power of two, standard techniques used in binomial trees are applied (e.g., forwarding from excess ranks (MPICH, 2025a)). Moreover, if the root is rank $t\neq 0$, we apply a logical rotation by subtracting $t$ from all rank identifiers modulo $p$.
2.3. Bine Trees Construction: Formal Definition
We now provide a formal construction process for distance-halving Bine trees. Table 1 summarizes the notation we use.
Table 1. Notation used.
Symbol Definition
$p$ Number of ranks.
$r$ Rank identifier.
$s$ Number of steps in a binomial or Bine tree collective ($s=\log_{2}p$).
$rank2nb(r)$ Convert rank identifier $r$ to its negabinary representation.
$nb2rank(r)$ Get rank identifier $r$ from its negabinary representation.
2.3.1. Ranks Representation
In Bine trees, each rank is assigned a negabinary (i.e., negative-base) representation, where numbers are expressed as sums of powers of 2-2 instead of $2$, as in standard binary. For example, the number $2$ is represented as $110_{-2}$, since $1\cdot(-2)^{2}+1\cdot(-2)^{1}+0\cdot(-2)^{0}=4-2=2$. Unlike standard binary, negabinary representations can encode both positive and negative integers (e.g., $011_{-2}=-1$). As a result, there is no one-to-one correspondence between non-negative ranks and their negabinary representation. The largest positive number $m$ that can be represented with a fixed number of negabinary bits is obtained by setting ones in all even positions and zeros in all odd positions. This ensures that only even powers of 2-2, which contribute positive values, are included. For instance, on six bits $m=010101_{-2}=16+4+1=21$.
The construction of Bine trees ensures that the rightmost rank is $m$. For example, for a 8-nodes tree we need three bits, and $m=101_{-2}=5$ (Fig. 3, [Uncaptioned image]). Therefore, all ranks in the range $[0,m]$ (to the right of rank 0) can be represented using their positive negabinary value (e.g., Fig. 3, [Uncaptioned image]). Ranks greater than $m$ (to the left of rank 0) are represented using the negabinary representation of rpr-p. For instance, rank $r=6$ in an 8-node tree is represented as $6-8=-2=010_{-2}$ (Fig. 3, [Uncaptioned image]).
We define the functions rank2nb(r,p){rank2nb}(r,p) and nb2rank(r,p){nb2rank}(r,p) to convert a rank identifier to its negabinary representation, and vice versa, for a collective on $p$ ranks. For example, ${rank2nb}(2,8)=110_{-2}$ and ${rank2nb}(6,8)=010_{-2}$. Both functions can be efficiently implemented using bit masking and an addition or subtraction. When the context is clear, we will omit the $p$ parameter and the -2 base.
2.3.2. Determining the Communication Partner
In the broadcast, each rank needs to determine when to join the tree, or in other words, at which step it will receive data from its parent. We denote by $u$ the number of consecutive least significant bits that are equal to each other, starting from the least significant bit. For instance, for a 16-node Bine tree, $u=3$ for $1000$, and $u=2$ for $1011$. In a distance-halving Bine tree broadcast, rank $r$ receives data at step i=sui=s-u (with $i\in[0,s-1]$). As illustrated in Fig. 4 [Uncaptioned image], rank $8$ receives the data from its parent at step $i=1$, since $rank2nb(8)=1000$, and thus $u=3$, giving i=su=43=1i=s-u=4-3=1.
Figure 4. A 16-node (order 4) distance-halving Bine tree.
[…]
A 16-node (order 4) distance-halving Bine tree.
Once a rank receives the data, in each subsequent step $i$ it forwards the data to rank:
(1) $q={nb2rank}({rank2nb}(r)\oplus\overbrace{111\ldots 1}^{s-i\>\text{bits}})$
with $\oplus$ denoting the XOR operation. In other words, at step $i$, each rank $r$ that has already received the data will send it to a rank $q$ whose negabinary representation differs in the least significant sis-i bits. For example, as shown in Fig. 4 [Uncaptioned image], at step $i=2$, rank $8$ sends data at rank $7$ since their negabinary representations ($1000$ and $1011$, respectively), differ for the least significant si=42s-i=4-2 bits. For clarity, the complete MPI code of the distance-halving Bine tree algorithm for the broadcast is provided in Listing 1. Notably, the implementation is no more complex than that of standard binomial trees (MPICH, 2025a; MPI, 2025a). For space reasons, we omit handling cases where the number of nodes is not a power of two; these can be managed similarly to standard binomial tree algorithms.
Listing 1: Bine tree distance-halving broadcast code.
Correctness can be easily demonstrated. For example, rank $4$ receives the data via $0\to 3\to 4$, corresponding to the XOR combination $0000\oplus 0111\oplus 0011$. More generally, any $s$-bit number can be uniquely represented as the XOR of a subset of the binary/negabinary numbers $\{1,11,111,\ldots,\underbrace{111\ldots 1}_{s\>\text{bits}}\}$, ensuring correctness.
2.3.3. Determining Ranks in a Subtree
As we will discuss in Sec. 4, collectives like scatter (Sec. 4.2) require identifying all ranks in the subtree rooted at a given rank. After receiving data at step $i$, a rank applies XOR on its least significant bits to determine its next destination. This process recurses, and at each step, the $i+1$ most significant bits remain unchanged. As a result, all descendants of a rank share the same $i+1$ leading bits in their negabinary representation. For instance, in a 16-node Bine tree, rank $8$ is reached at step $i=1$, so all the ranks in the subtree rooted in $8$ share its two most significant bits (i.e., $10$, as shown in Fig. 4, [Uncaptioned image]).
2.4. Advantages over Binomial Trees
We know quantify the maximum reduction in global links traffic both theoretically (Sec. 2.4.1) and empirically (Sec. 2.4.2).
2.4.1. Theoretical Bound Computation
To explain why we expect Bine trees to reduce the number of bytes on global links compared to standard binomial trees, we start by considering the distance between communicating ranks. In standard distance-halving binomial trees, two communicating ranks differ for the bit in position si1s-i-1 (i.e., the distance is $\delta_{{binomial}}(i)=2^{s-i-1}$). In distance-halving Bine trees [^3] Although we refer to it as a distance-halving Bine tree for simplicity, the distances do not halve exactly; instead, they differ by at most $\pm 1$ from the ideal halving., however, the negabinary representations of the two communicating ranks differ in the sis-i least significant bits, which are all equal to 11 for one rank, and all equal to 0 for the other rank. Thus, the distance between the two ranks at step $i$ is: $\delta_{bine}(i)=\lvert\sum_{j=0}^{s-i-1}(-2)^{j}\rvert=\lvert\frac{1}{3}-\frac{1}{3}(-2)^{s-i}\rvert\approx\frac{2^{s-i}}{3}$.
We can estimate the reduction in the distance between communicating ranks by considering the ratio:
(2) $\frac{\delta_{bine}(i)}{\delta_{binomial}(i)}=\frac{2^{s-i}}{3\cdot 2^{s-i-1}}=\frac{2}{3}$
This implies that in Bine trees, communicating ranks are at a $\sim 33\%$ shorter distance compared to standard binomial trees. As a result, the probability that two communicating ranks are in different groups and thus need to communicate through global links is reduced. Quantifying the exact reduction in inter-group traffic is challenging, as it heavily depends on the specific topology and on how ranks are allocated to groups. However, we can estimate an upper bound by considering a worst-case scenario in which each group only has a single outgoing link to another group.
Let us focus on a specific step $i$ and assume the group size equals $\delta_{binomial}(i)$, such that all communications in the standard binomial tree cross group boundaries. In this worst-case scenario, all $\delta_{binomial}(i)$ communications would be directed to nodes in other groups, whereas for Bine trees this would only happen for $\delta_{bine}(i)$ communications. Therefore, the maximum reduction in global link traffic corresponds to the ratio $\delta_{bine}(i)/\delta_{binomial}(i)$, which, as shown in Eq. 2, is equal to $2/3$. In other words, Bine reduces by at most $33\%$ the traffic on global links.
2.4.2. Empirical Analysis
We additionally validated our assumptions empirically by analyzing one-week job allocation data from the Leonardo (Turisini et al., 2024) supercomputer and two-weeks data from the LUMI (EuroHPC Joint Undertaking, [n. d.]) supercomputer. Using Slurm commands (squeue and scontrol), we obtained the list of jobs running on the system and, for each job, the hostnames of the allocated nodes. We then identified the Dragonfly/Dragonfly+ group to which each node was assigned. On Leonardo, we relied on existing hostname-to-group mappings (CINECA, 2025), while on LUMI we extracted this information using the sinfo -N -o "%N %f" command. With this data, we determined the group assignment of each rank for every job. Based on this mapping we estimated, for each job, the global traffic of an allreduce operation using standard binomial and Bine trees, and computed the resulting traffic reduction achieved by Bine across allocations.
Figure 5. Reduction in global traffic on Leonardo and LUMI when using Bine trees compared to standard binomial trees.
[…]
Reduction in global traffic on Leonardo and LUMI when using Bine trees compared to standard binomial trees.
We present this analysis in Fig. 5, where we report the distribution of global traffic reduction across jobs using boxplots, grouped by the number of nodes. The data includes a total of $1116$ jobs on Leonardo and $1914$ jobs on LUMI. The largest observed jobs spanned up to $256$ nodes on Leonardo and $2048$ nodes on LUMI, matching the maximum job sizes allowed on the two systems. For clarity, outliers are not shown in the figure. No outliers above the theoretical upper bound of 33% were observed, coherently with our model (Eq. 2). A few outliers reached values as low as $200\%$, all occurring in small allocations (fewer than 64 nodes). As discussed in Sec.2.2, this happens because Bine minimizes distance modulo $p$ rather than the actual distance. In a few small allocations, this can increase global traffic compared to binomial trees, although on average, they remain comparable. In contrast, for larger allocations, Bine consistently outperforms standard binomial trees. The reduction in global link traffic increases with the number of nodes, as larger jobs typically span more groups, creating more opportunities for Bine to reduce inter-group communication.
3. Bine Butterflies and Distance-Doubling Trees
3.1. Bine Butterflies
In a butterfly communication pattern, each rank exchanges data with another rank at every step. A Bine distance-halving butterfly can be seen as the superposition of multiple distance-halving Bine trees, each rooted at a different rank. The left part of figure 6 illustrates this view, highlighting the trees rooted at ranks 2 and 3 with solid orange and dotted red lines, respectively. To guarantee correctness, these trees must be arranged so that the send and receive decisions are consistent across all ranks.
Figure 6. An 8-node distance-halving Bine butterfly (left), and distance-doubling Bine butterfly (right).
[…]
An 8-node distance-halving Bine butterfly (left), and distance-doubling Bine butterfly (right).
We take as a reference the distance-halving tree rooted in rank 0. As discussed in Sec. 2.3, at step $i$ rank 0 communicates with a rank whose negabinary representation has the least significant sis-i bits set to 1, and the remaining bits set to 0, i.e., with rank:
(3) $q=\Bigg{(}\sum_{j=0}^{s-i-1}(-2)^{j}\Bigg{)}\bmod p=\Big{(}\frac{1}{3}-\frac{(-2)^{s-i}}{3}\Big{)}\bmod p$
Trees rooted in even ranks $r$ have the same structure as the tree rooted in 0, but rotated to the right by $r$ positions. Moreover, since each rank communicates with another rank whose negabinary representation differs in the last $i+1$ bits, even ranks always communicate with odd ranks, and vice versa. Therefore, the tree rooted at an odd rank $r$ must be mirrored with respect to the tree rooted at the even rank r1r-1. Thus, we can determine the communication partner $q$ for any rank $r$ at step $i$ as:
(4) $q=\begin{cases}\Big{(}r+\frac{1}{3}-\frac{(-2)^{s-i}}{3}\Big{)}\bmod p&\text{if }r\text{ is even}\\ \Big{(}r-\frac{1}{3}+\frac{(-2)^{s-i}}{3}\Big{)}\bmod p&\text{if }r\text{ is odd}\end{cases}$
Since the distance between communicating ranks in Bine butterflies matches that of Bine trees, distance-halving Bine butterflies reduce global traffic by up to $33\%$ compared to standard distance-halving butterflies. Distance-halving butterflies are advantageous for collectives that increase the size of the exchanged data at each step, such as allgather, since in the last steps, where more data is transmitted, communication happens between closer nodes. Alternatively, it is possible to build a distance-doubling Bine butterfly by inverting the communication order, as shown in the right part of Fig. 6. This is preferable for collectives that exchange more data in the first steps, such as reduce-scatter (Sec. 4.3).
3.2. Distance-Doubling Bine Trees
As with the distance-halving Bine tree, we first describe the case where the root is at rank 0. For a tree rooted at any other rank $t$, we follow the same approach as in distance-halving trees, by rotating rank identifiers to the right by $t$ positions.
3.2.1. Ranks Representation
We assign to each rank $r$ a value $h(r,p)$. For even ranks, h(r,p)=rank2nb(pr)h(r,p)=rank2nb(p-r), whereas for odd ranks $h(r,p)=rank2nb(r)$. Then, we assign to each rank $r$ a value $\nu(r,p)=h(r,p)\oplus(h(r,p)>>1)$ (>>>> is the logical right shift). Examples for even and odd ranks are shown in the right part of Fig. 6. When clear from the context, we omit $p$ for the sake of clarity.
3.2.2. Determining the Communication Partner
The communication partner of a rank $r$ at step $i$ is the rank $q$ such that $\nu(q)=\nu(r)\oplus 2^{i}$ (i.e., a rank that differs in the $i$-th bit in the $\nu$ representation). A rank communicates only after it receives the data from its parent. This happens at step $i$, where $i$ is the position of the most significant bit set to 11 in $\nu(r)$. In the example in the right part of Fig. 6, rank $2$ receives the data from its parent at step $1$, since the most significant bit set to 11 in $\nu(2)=011$ is in position $1$. Then, at step $2$, it sends the data to a rank whose $q$ such that $\nu(q)=011\oplus 2^{2}=011\oplus 100=111$ (i.e., rank $5$). The algorithm is correct because it operates in the same manner as the standard binomial tree algorithm, but using $\nu(r)$ instead of $r$.
3.2.3. Determining Ranks in a Subtree
All ranks in a subtree rooted at $r$ share the same $i+1$ least significant bits in their $\nu$ representation, where $i$ is the step at which $r$ received the data from its parent. Indeed, after receiving the data at step $i$, rank $r$ applies XOR on the bits in positions greater than $i$. This process recurses, and at each step the $i+1$ least significant bits remain unchanged. For example, in the right part of Fig. 6, rank $1$ receives the data at step $i=0$, and its $i+1=1$ least significant bit in $\nu(1)$ set. Consequently, all its descendants are ranks $q$ whose $\nu(q)$ also have the least significant bit set. It is worth noting that, unlike distance-halving Bine trees, nodes in distance-doubling Bine subtrees are not contiguous. For some collective operations, this may require transmitting non-contiguous data (see Sec. 4.3.1).
4. Collectives Design
4.1. Gather
Similarly to standard binomial trees, in a Bine gather each rank gathers data from another rank at each step, halving the number of active ranks until only the root remains. We illustrate in Fig. 7, a standard binomial gather and a Bine gather over eight ranks, both using distance-halving trees. We consider data to be divided in “blocks”, one per rank participating in the gather. Let $[a,b]$ denote the range of blocks a rank holds at any time. When sending, a rank transmits all the blocks it currently has. In a standard binomial gather, if a rank holds $k$ blocks, it always receives the next $k$ contiguous blocks: that is, $[b+1,b+k]$. E.g., at step 1, rank 4 has blocks $[4,5]$ and receives $[6,7]$.
Figure 7. 8-node gather with binomial and Bine trees.
[…]
An 8-node gather with standard binomial trees and Bine trees.
In Bine trees, however, the range may extend in either direction. A rank holding $[a,b]$ might receive either $(b+1)\bmod p,(b+k)\bmod p]$ or $(a-k)\bmod p,(a-1)\bmod p]$, extending its buffer upward or downward. For example, at step 1, rank 0 with blocks $[0,1]$ receives $[6,7]$ from rank 7, extending its buffer upward. In general, even ranks start extending downward, odd ranks upward, alternating direction at each step (e.g., rank 0 extends downward at step 0, upward at step 1, and downward again at step 2).
4.2. Scatter
For the scatter we can use the opposite process of that used for the gather. In this case, at each step, each rank sends either the bottom or the top half of its remaining buffer $[a,b]$. I.e., if a rank has $k$ blocks of data, it can either send the blocks $a,(a+\frac{k}{2})\bmod p]$, or the blocks $(b-\frac{k}{2})\bmod p,b]$ For example, since in the gather rank 0 received in the last step the sub-buffer with range $[2,5]$, then it means that in the scatter it needs to send the sub-buffer with range $[2,5]$ in the first step. Thus, we need to know what would have been the values of $[a,b]$ in the last step of the gather.
In the gather, even ranks extended their buffer by adding $2^{0}+2^{2}+2^{4}+\ldots$ to $b$ and subtracting $2^{1}+2^{3}+2^{5}+\ldots$ from $a$. Odd ranks did the opposite: they subtracted $2^{0}+2^{2}+2^{4}+\ldots$ from $a$ and added $2^{1}+2^{3}+2^{5}+\ldots$ to $b$. This corresponds to adding or subtracting the binary numbers $0101\ldots 0101_{2}$ or $1010\ldots 1010_{2}$ to $a$ or $b$. Thus, it is enough to add and subtract those quantities to the rank identifier $r$ to find the starting $[a,b]$ for the scatter, depending on whether the rank is even or odd. For example, in the first step of the scatter, rank 0 has $[a,b]=[6,5]$. Then, at each step, either $b$ is decreased by $\frac{k}{2}$ or $a$ is increased by $\frac{k}{2}$, depending on whether the rank is even or odd. This process continues until $k=1$.
4.3. Reduce-Scatter and Allgather
We discuss in this section the design of the reduce-scatter collective. For the allgather, it is enough to reverse the reduce-scatter communication pattern. We consider a reduce-scatter algorithm based on a vector-halving, distance-doubling butterfly pattern, using the Bine butterfly introduced in Sec. 3.1. In this scheme, each rank sends half of its data to a peer and receives half of the peer’s data at each step. Since the communication volume is highest in the early steps, relying on a distance-doubling butterfly helps reduce global traffic. As in standard recursive halving reduce-scatter, each rank sends a total of $n\frac{p-1}{p}$ bytes over $\log_{2}p$ steps.
A distance-doubling butterfly reduce-scatter can be conceptually viewed as the concurrent execution of a distance-doubling scatter and a distance-halving reduce. Focusing on the scatter component, unlike the scatter described in Sec. 4.2, a distance-doubling tree is used. While this reduces the volume of inter-group communication, it complicates data selection: a rank $r$ must send to a rank $q$ all blocks associated with the ranks in the subtree rooted at $q$. As discussed in Sec. 3.2.3, these subtrees consist of non-contiguous ranks, requiring the exchange of non-contiguous blocks.
4.3.1. Dealing with Non-Contiguous Data
Non-contiguous transmissions can hinder performance, especially for small- to medium-sized vectors (Loch and Koslovski, 2021). This can be mitigated in several ways. One option is to use MPI datatypes to handle non-contiguous data or to copy the data to a temporary buffer before transmission. However, this might still introduce a considerable overhead.
Alternatively, when $p$ is a power of two, we can permute the data upfront, ensuring subsequent transmissions use contiguous data. This can be achieved by moving the data blocks from position $r$ to position ${reverse}(\nu(r))$, where reverse{reverse} reverses the order of the bits in $\nu(r)$. As discussed in Sec. 3.2.3, all descendants of a given rank share a fixed number of least significant bits in their $\nu$ representation. By reversing $\nu(r)$, all blocks sent by rank $r$ would then share a fixed number of most significant bits, and therefore be contiguous. Another option is to skip the permutation and have ranks send contiguous data as if the permutation had been applied. For the reduce-scatter, this means that at the end of the collective each rank would hold the block of another rank. A final communication step is then sufficient to exchange data and reorder it correctly. Alternatively, some collectives combine two operations (e.g., an allreduce as a reduce-scatter followed by an allgather). In such cases, even if ranks do not receive the correct block after the first operation, the subsequent collective (e.g., the allgather) reverses the permutation and reorders the blocks correctly.
4.4. Allreduce and Alltoall
For the allreduce, we use two algorithms: one for small and one for large vectors. As for standard binomial tree algorithms, the former performs fewer steps, whereas the latter performs more steps but transfers less data. For small vectors, we employ a standard recursive doubling allreduce using Bine butterflies. For large vectors, we perform a Bine reduce-scatter followed by a Bine allgather. Both algorithms create a communication pattern similar to the Swing algorithm (De Sensi et al., 2024a). However, unlike Swing, which always sends non-contiguous data (Daniele De Sensi, 2025) and needs to issue multiple sends at each step, Bine sends contiguous data (when $p$ is a power of two), leading to $2\times$ performance improvements (see Sec. 5.2.2).
Conceptually, the alltoall can be viewed as a small vector allreduce, where ranks send at each step $n/2$ bytes, and the received data is concatenated with their own rather than being aggregated. At each step, each rank moves the data it wants to keep to the left of its buffer and the data it needs to send to the right, similar to the rotations in Bruck’s algorithm. Each rank tracks the block stored in each buffer position. At the end of the algorithm, a final permutation step is needed to put the data in the correct order.
4.5. Reduce and Broadcast
For Bine reduce and broadcast, we propose algorithms for both small and large vectors. The small vector algorithm uses distance-halving Bine trees for broadcasting data from the root to the leaves or reducing data from the leaves to the root. For large vectors, we implement the reduce as a reduce-scatter followed by a gather, and the broadcast as a scatter followed by an allgather. The reduce uses a distance-doubling Bine butterfly for the reduce-scatter and a distance-halving Bine tree for the gather. Data is transmitted contiguously, without the need for reordering, as the gather inverts the block permutation done by the reduce-scatter, placing the blocks back in the correct order. Similarly, the broadcast relies on distance-doubling scatter and distance-halving allgather.
5. Experimental Evaluation
We evaluate Bine trees on four supercomputers ranked among the top 11 in the Top500 list at the time of writing. Table 2 summarizes their network topologies and the MPI libraries used, as recommended by their respective supercomputing centers. Additional system-specific details appear in later sections. Our selection covers all four commonly used oversubscribed network topologies, providing a broad and representative evaluation.
Table 2. Systems used in the experimental evaluation. Top500 rankings taken from November 2024 list.
System (Top500 Rank) Topology MPI Library
LUMI (#8) Dragonfly Cray MPICH v8.1.29
Leonardo (#9) Dragonfly+ Open MPI v4.1.6
MareNostrum 5 (#11) 2:1 Oversubscribed Fat Tree Open MPI v4.1.5
Fugaku (#6) 6D Torus Fujitsu MPI v4.0.1
Implementation
We implemented all Bine algorithms from Sec. 4 on top of MPI, without modifying any MPI library. This choice was motivated by both portability (our target systems rely on four different MPI libraries, see Table 2) and fairness, allowing for direct comparison against state-of-the-art algorithms using the same MPI backend. To validate this design choice, we implemented Bine versions of allreduce, allgather, and reduce-scatter within Open MPI and compared them to their counterparts built on top of MPI. The observed performance difference was below 3%. Therefore, in the following, we report results using the Bine implementations built on top of MPI. This was not the case on Fugaku, where, to concurrently use all the available NICs, our implementation uses the low-level uTofu library (Fujitsu, 2025) (see Sec. 5.4). We implemented all Bine algorithms in the PICO benchmarking framework (of Rome, 2025b), which streamlines comparison with the collective algorithms provided by default in Open MPI and MPICH. Additionally, we provide self-contained reference implementations (of Rome, 2025a).
Algorithms Used in the Comparison
We compare Bine trees against all the collective algorithms available in the MPI implementations on each system. To ensure a fair comparison, we evaluate Bine both against the default algorithm selected by MPI and each available algorithm individually, by manually enforcing the selection. We also added other state-of-the-art algorithms from other MPI libraries or versions, such as the sparbit allgather algorithm (Loch and Koslovski, 2021; MPI, 2025d) (not available on Open MPI v4.1.x), and the ring-based allreduce algorithm (not available in MPICH 3.4a2, on which Cray MPICH v8.1.29 is based (MPICH, 2025)). Moreover, we implemented the Swing algorithm (De Sensi et al., 2024a) for allgather, reduce-scatter, and allreduce, and the torus-optimized Bucket algorithm (Jain and Sabharwal, 2010; Barnett et al., 1995) (essentially a multi-dimensional ring) for the same collectives. For each system, we begin with a focused comparison of Bine and binomial trees, and then extend the evaluation to include all available algorithms.
Benchmarking Methodology
For each collective, we measured performance across different node counts (without requesting any specific node placement) and vector sizes ranging from 32 B to 512 MiB. Due to space constraints, we report results only for power-of-two node counts. Results for non-power-of-two cases are consistent with those shown and are available in the reproducibility package. All collectives operate on vectors of 32-bit integers. For each combination of node count and vector size, we executed each collective up to $20000$ times for small vectors and five times for 512 MiB vectors. Most experiments were conducted with one process per node; performance with multiple processes per node is discussed in Sec.6. All vectors are located in host memory except for the experiments in Sec. 6 where we analyze multi-GPU collectives. We followed standard benchmarking practices (Hoefler and Belli, 2015), by discarding the first 20% of iterations as warm-up, and reporting for each iteration the maximum runtime across all ranks.
5.1. Evaluation on LUMI
LUMI (EuroHPC Joint Undertaking, [n. d.]) relies on a 24-group Slingshot 11 Dragonfly network (De Sensi et al., 2020), with 124 nodes per group. We used the LUMI-G partition, where each node has a 64-core AMD EPYC™ Trento 7A53 CPU (organized into four NUMA domains with 128GB of DDR4 memory each), four 200Gb/s Cassini-1 NICs, and four AMD MI250x GPUs.
5.1.1. Comparison with Binomial Trees
We compare Bine and binomial trees in terms of both performance and the reduction in the number of bytes transmitted over global links. To assess this, we calculated the total bytes exchanged between ranks. By using the node locations from the /etc/cray/xname file, we identified the group each node belongs to and quantified the traffic forwarded between ranks assigned to different groups.
On LUMI, we conducted experiments using between $16$ and $1024$ nodes (the maximum allowed) spanning from one to 21 groups. As discussed in Sec. 4.3, Bine ’s reduce-scatter and allreduce algorithms may send multiple non-contiguous data blocks, rather than a single bulk transfer. This design improves the overlap between communication and computation in reduce-scatter. In contrast, the standard MPICH binomial tree and butterfly implementations transmit large contiguous blocks, without exploiting such overlap. To ensure a fair comparison focused on algorithmic structure rather than implementation-level optimizations, we excluded these Bine variants from the analysis. For the alltoall collective, we compared Bine with Bruck’s algorithm, as it is the closest to binomial algorithms, performing a number of steps logarithmic in $p$.
Given the large volume of data collected, across eight collectives, up to ten algorithms per collective, nine vector sizes, and up to seven node counts, for a total of over $5000$ configurations, we present a summarized analysis in Table 3444The complete set of detailed results is available in the reproducibility package.. For each collective, we report the fraction of configurations (i.e., combinations of node count and vector size) in which Bine outperforms binomial trees and vice versa (“%Win/%Loss”). When Bine is superior, we report both the average and maximum performance improvements over all the configurations. We compute the average using the geometric mean, which is more appropriate for summarizing ratios (Hoefler and Belli, 2015). Similarly, when binomial trees outperform Bine, we report the corresponding average and maximum performance degradations.
Lastly, we report both the average and maximum reduction in global-link traffic. Here, global links refer to those connecting switches in different Dragonfly groups. We assume packets traverse inter-group connections via minimal paths, an assumption that may not always hold in practice, especially in low-diameter networks such as Dragonfly (De Sensi et al., 2020). As a result, the reductions we report should be interpreted as lower bounds; the traffic savings induced by Bine trees could be even greater than our estimation.
Table 3. Comparison with Binomial Trees on LUMI
Coll. % Win Avg/Max Perf. Gain % Loss Avg/Max Perf. Drop Avg/Max Traffic Red.
Allreduce 67% 7%/67% 14% 9%/13% 11%/20%
Allgather 47% 19%/177% 41% 11%/17% 11%/20%
Red.-Scat. 39% 21%/143% 29% 13%/29% 11%/20%
Alltoall 94% 33%/191% 2% 16%/16% 15%/20%
Bcast 67% 25%/79% 27% 16%/29% 88%/94%
Reduce 87% 7%/32% 5% 10%/14% 10%/19%
Gather 71% 13%/47% 16% 13%/32% 9%/25%
Scatter 61% 7%/30% 24% 9%/20% 9%/25%
Overall, Bine outperforms standard binomial trees in more than 60% of the configurations on most collectives. The only exceptions are reduce-scatter and allgather, where Bine still performs better more often than it is outperformed by binomial trees. For allgather, Bine is superior in 47% of the cases, while binomial trees outperform it in 41%, with the remaining 12% showing minimal differences (below 1%). The smaller advantage for these two collectives is due to the additional data permutation step required by Bine trees, either through local buffer shuffling or additional data communication (as discussed in Sec. 4.3). However, when Bine outperforms binomial trees, the average performance gain is around 20%, with a maximum improvement of up to 177%.
Notably, for alltoall, Bine outperforms Bruck’s algorithm in 94% of the cases. Given that alltoall typically involves large data transfers, reducing the traffic on global links becomes especially important. However, as we will discuss in Sec. 5.1.2, logarithmic alltoall algorithms like Bruck’s and Bine are typically only effective for small messages and large node count, while linear algorithms tend to perform better in the other cases. It is worth remarking that, as is typical for collective operations, the optimal algorithm depends on both the vector size and the number of nodes. As a result, no single algorithm consistently outperforms the others across all scenarios, and such trade-offs are therefore expected for all collectives.
Bine reduces traffic on global links by an average of 10% across all collectives, and up to 20% at larger node counts, as more nodes lead to higher utilization of global links. The largest reduction, up to 94%, is observed in broadcast. This is due to MPICH (and Open MPI, as shown later) using a large-vector broadcast based on a distance-halving scatter, followed by a distance-doubling allgather (MPICH, 2025b; MPI, 2025c). In contrast, Bine trees reverse this process, significantly reducing traffic on global links, particularly in the allgather phase.
5.1.2. Comparison with Other State-of-the-Art Algorithms
To provide a complete comparison, we now compare Bine with state-of-the-art algorithms (including binomial trees as well).
Allreduce
We present a detailed analysis of the allreduce in Fig. 8(a), using a heatmap to display all combinations of node count and vector size. Each cell either indicates the best-performing algorithm with a single letter or, if Bine is superior, shows the performance ratio between Bine and the best state-of-the-art algorithm. Bine outperforms other algorithms in almost all configurations, with peak performance gains of up to $1.62\times$ on medium-sized vectors and high node counts.
Other Collectives
For the remaining collectives, we summarize the corresponding heatmaps using boxplots in Fig. 8(b), retaining only the configurations where Bine outperforms all other algorithms. Below each collective name, we report the percentage of such cases. For reduce-scatter, we observe up to 65% performance improvements for medium-sized vectors. As discussed in Sec. 4.3 reduce-scatter either sends non-contiguous blocks or permutes the buffer and sends contiguous blocks. At large node counts, permuting the vector for contiguous data results in a 20% performance drop for Bine compared to binomial trees, as discussed in Sec. 5.1.1.
For allgather, Bine was outperformed by the default Cray MPICH algorithm in 17% of the cases. According to the Cray MPICH manual, the default algorithm leverages architecture-specific optimizations, which we suppose may include, for example, triggered operations. By disabling such optimizations, Bine outperformed the state of the art in an additional 18% of cases, with maximum performance gains reaching 89%. A similar trend was observed for broadcast, with Bine becoming the top-performing algorithm in 63% of cases when disabling architecture-specific optimizations. For alltoall, Bine is the best algorithm in 21% of cases. On 512 and $1024$ nodes with 128 KiB vectors, it outperforms all others by up to 78%. Like Bruck’s algorithm, Bine is optimized for small vectors, favoring fewer communication steps at the cost of more data per step. Hence, it performs best at larger scales and with smaller vectors.
(a) Allreduce
(b) All collectives
Figure 8. Comparison with state-of-the-art algorithms on LUMI. (a): Each cell shows a letter indicating the best-performing algorithm (N = binomial, D = default), or, when Bine is best, the performance ratio between Bine and the next-best algorithm. (b): For each collective, the number below the name indicates the percentage of tests where Bine outperforms binomial trees. Box plots summarize the distribution of performance improvements in those cases.
[…]
Comparison of Bine performance against state-of-the-art algorithms on LUMI. Left: Each cell shows either a letter indicating the best-performing algorithm (N = binomial, D = default), or, when Bine is best, the performance ratio between Bine and the next-best algorithm. Right: For each collective, the number below the name indicates the percentage of configurations where Bine was the top performer. Box plots summarize the distribution of performance gains in those cases.
5.2. Evaluation on Leonardo
Leonardo (Turisini et al., 2024) uses a Dragonfly+ network topology based on Infiniband HDR. It consists of 23 groups, each containing 180 nodes, organized as two-level fat trees. We used the Booster partition, where each node is equipped with two 200 Gb/s dual-port NVIDIA Connect-X6 NICs, a single-socket 32-core Intel Xeon®8358 processor, 512 GB of host memory, and four NVIDIA A100 GPUs.
Table 4. Comparison with Binomial Trees on Leonardo
Coll. % Win Avg/Max Perf. Gain % Loss Avg/Max Perf. Drop Avg/Max Traffic Red.
Allreduce 67% 11%/46% 20% 10%/13% 19%/26%
Allgather 91% 9%/54% 0% 0%/0% 19%/26%
Red.-Scat. 71% 4%/12% 14% 6%/6% 17%/23%
Alltoall 79% 25%/70% 0% 0%/0% 15%/15%
Bcast 94% 41%/148% 0% 0%/0% 89%/92%
Reduce 44% 43%/72% 33% 7%/9% 13%/19%
Gather 93% 23%/71% 7% 12%/12% 12%/20%
Scatter 93% 22%/63% 0% 0%/0% 12%/20%
5.2.1. Comparison with Binomial Trees
Table 4 compares Bine to binomial trees on configurations from $16$ to $2048$ nodes, spanning 3 to 20 groups. Since standard users on Leonardo are limited to jobs of at most $256$ nodes, the experiments on $512$, $1024$, and $2048$ nodes were conducted during a maintenance window, with support from CINECA, on an otherwise idle system. To reduce the test duration and minimize impact on system operations, we collected data for only allreduce and allgather on jobs larger than $256$ nodes.
The results show that Bine outperforms binomial trees in most scenarios. For half of the collectives, Bine is the best-performing algorithm in over 90% of the cases and matches the performance of binomial trees in the remaining ones.. On broadcast we see higher advantages than on LUMI, especially for vectors smaller than 128 KiB, where performance gain where of 4% on LUMI but of 42% on Leonardo. This can be attributed to MPICH relying on distance-halving broadcast (MPICH, 2025a), whereas Open MPI relies on distance-doubling broadcast (MPI, 2025a, b), which increases the number of bytes transmitted over global links, as we shown in Fig. 1.
5.2.2. Comparison with Other State-of-the-Art Algorithms
Allreduce
We compare allreduce performance with state-of-the-art algorithms in Fig. 9(a). Bine emerges as the best-performing algorithm in 67% of the configurations, achieving performance improvements of up to $1.45\times$. For large vectors at small node counts, the ring algorithm is usually more effective. We verified that part of Bine ’s performance gains can be attributed to its use of segmentation during allreduce, which enables better overlap between communication and computation. To isolate segmentation effects, we implemented a version of Bine allreduce without segmentation.
Without segmentation, the ring algorithm outperforms Bine for 512 MiB messages on $256$ and $512$ nodes. This is expected, as the ring algorithm inherently splits the data into smaller chunks than both Bine and butterfly algorithms, which enhances overlap. To have a fair comparison with the ring algorithm, we present Bine’s results with segmentation enabled. Still, except for these two cases, Bine outperforms all other algorithms even without segmentation.
Other Collectives
We summarize the results in Fig. 9(b). For allgather on $2048$ nodes, vectors smaller than 64 MiB benefit from the remapping approach described in Sec.4.4, achieving $2\times$ performance improvements compared to algorithms like Swing (De Sensi et al., 2024a) or sparbit (Loch and Koslovski, 2021; MPI, 2025d), which send non-contiguous data and introduce significant overhead. For reduce-scatter, Bine is the top-performing algorithm in 63% of the cases. For alltoall, results are similar to those on LUMI, with Bine being the best algorithm on small vectors. Finally, Bine achieves higher performance on broadcast compared to LUMI, consistently with our findings in Sec. 5.2.1.
(a) Allreduce
(b) All collectives
Figure 9. Comparison with state-of-the-art algorithms on Leonardo (N = binomial, D = default, R = ring).
[…]
Comparison with state-of-the-art algorithms on Leonardo (N = binomial, D = default, R = ring).
5.3. Evaluation on MareNostrum 5
MareNostrum 5 (Banchelli et al., 2025) uses a 2:1 oversubscribed fat-tree network topology based on InfiniBand NDR200. Each full-bandwidth sub-tree consists of 160 nodes. We utilized the ACC partition, where each node is equipped with $2\times$ Intel Sapphire Rapids 8460Y+ CPUs, 512 GB of DDR5 memory, and four 200 Gb/s NICs. Additionally, each node has 4x Nvidia Hopper GPUs with 64 GB of HBM2 memory.
5.3.1. Comparison with Binomial Trees
We ran experiments from $4$ to $64$ nodes (the maximum allowed), spanning between one and eight subtrees. We report the results of our analysis in Table 5. Consistent with our findings on LUMI and Leonardo, Bine trees outperform binomial tree algorithms in most cases, with performance improvements of up to 158%. However, in a few cases (more evident for gather and scatter), Bine increases (on average) the traffic on global links compared to binomial trees. We observed this mostly for small node count, consistently with what we discussed in Sec. 2.
Table 5. Comparison with Binomial Trees on MareNostrum 5
Coll. % Win Avg/Max Perf. Gain % Loss Avg/Max Perf. Drop Avg/Max Traffic Red.
Allreduce 80% 9%/46% 2% 6%/6% 2%/11%
Allgather 55% 10%/42% 24% 12%/17% 0%/11%
Red.-Scat. 74% 15%/78% 2% 6%/6% 0%/11%
Alltoall 88% 26%/62% 10% 11%/12% 19%/33%
Bcast 98% 32%/158% 0% 0%/0% 49%/86%
Reduce 51% 31%/80% 22% 19%/40% 2%/7%
Gather 69% 25%/76% 2% 8%/8% -8%/18%
Scatter 95% 36%/140% 5% 7%/8% -8%/18%
5.3.2. Comparison with Other State-of-the-Art Algorithms
Fig. 10(a) summarizes the comparison between Bine and the best state-of-the-art algorithms. The results are consistent with our observations on the other platforms. For alltoall, gather, and scatter, Bine is the best-performing algorithm in fewer cases. This is primarily due to the smaller scale of the experiments (up to 64 nodes), compared to $1024$ or $256$ nodes on LUMI and Leonardo. At these scales, linear algorithms often outperform logarithmic ones, especially for large vector sizes. Conversely, for vectors smaller than 1 MiB, Bine outperformed the other algorithms by up to $2\times$ on 64 nodes for scatter.
(a) MareNostrum 5
(b) Fugaku
Figure 10. Comparison with state-of-the-art algorithms.
[…]
Comparison with state-of-the-art algorithms.
5.4. Evaluation on Fugaku
Fugaku (Sato et al., 2022) is based on the Tofu Interconnect D (Ajima et al., 2018) and uses a 6D toroidal network. In practice, each job can be allocated a 3D sub-torus. The system is composed of almost 159,000 nodes, each equipped with a Fujitsu A64FX CPU, 32 GiB of HBM memory, and six TNIs (i.e., NICs), each capable of injecting 54.4 Gb/s.
5.4.1. Bine Optimization for Torus Networks
So far, we have not made any assumptions about the network topology. However, we can extend Bine trees to better exploit communication locality on torus networks. Instead of considering a flat rank space, we treat ranks as coordinates in a multidimensional space, reducing distance between communicating ranks, as typical in algorithms optimized for toroidal networks (Jain and Sabharwal, 2010; De Sensi et al., 2024a). Since each node has six NICs, we divide the vector used in the collective into six parts. Then, we execute six collectives in parallel, each operating on a different part of the vector, using a separate NIC, and transmitting in a different direction of the 6D torus. To ensure that each collective uses a different NIC, we implemented the Bine algorithms directly on top of the low-level uTofu library (Fujitsu, 2025).
5.4.2. Comparison with Binomial Trees
When comparing Bine to binomial trees, we observed up to $40\times$ speedups. However, this is largely because these binomial algorithms, based on Open MPI, on which Fujitsu MPI is based, are not optimized for toroidal networks.
5.4.3. Comparison with Other State-of-the-Art Algorithms
As with other systems, we compare both with the default algorithm selected by Fujitsu MPI and with all other available algorithms by explicitly forcing their selection. We performed our analysis on 2x2x2, 4x4x4, 8x8x8, 64x64, and 32x256 nodes jobs. We report results in Fig. 10(b). For allreduce, Bine trees outperformed state-of-the-art algorithms in 62% of cases, especially for vectors larger than 2 KiB. The highest gains—up to $4.12\times$—were seen for 128 KiB–64 MiB vectors on $8192$ (32x256) nodes. While Fugaku uses the torus-optimized Trinaryx algorithm (Ajima et al., 2018; Hosono et al., 2021; Fujitsu, 2025), which builds three edge-disjoint spanning trees, its linear number of steps limits scalability compared to Bine’s logarithmic approach.
For allreduce, reduce-scatter, and scatter, Bine trees were the top-performing algorithm in over 60% of the tests, with performance gains of up to $5\times$ for reduce-scatter. In contrast, for broadcast and reduce, Fujitsu MPI includes algorithms that are optimal on multiported torus networks, leaving little room for improvement. Nonetheless, the fact that Bine remains competitive, even against these highly optimized vendor implementations, underscores its strong generalization capability and broad effectiveness.
It is worth noting that Fugaku exhibits the highest performance gains among the four analyzed systems. These gains stem from Bine’s ability to reduce traffic on oversubscribed links. In the other three topologies, oversubscription primarily affects global links. In contrast, on a torus, all links can be considered oversubscribed. From another perspective, torus networks have the lowest bisection bandwidth among the topologies considered, and thus benefit the most from Bine’s enhanced locality.
6. Discussion
6.1. Impact of Processes per Node
All experiments in Sec. 5 were conducted with one process per node. To study the impact of multiple processes per node, we ran all collectives on 64 nodes on LUMI with either one or four processes per node (LUMI has four NICs per node). Performance was largely consistent, though in a few cases Bine achieved larger gains with four processes per node; for example, the 1 MiB reduce-scatter gain increased from 59% to 84%. This is likely due to the higher traffic generated by each node when using multiple processes, which better highlights Bine ’s global traffic reduction benefits. We observed similar trends on Leonardo and MareNostrum 5.
6.2. Multi-GPU Evaluation
We implemented a GPU-aware version of the Bine allreduce algorithm on top of MPI. This was motivated by two main goals. First, it allows us to evaluate the performance of Bine on multi-GPU systems and directly compare it against NCCL’s algorithms. Second, it enables the evaluation of a hierarchical implementation of Bine. We optimized our design and implementation for large vectors (i.e., implementing the allreduce as a reduce-scatter followed by an allgather), and for the MareNostrum 5 and Leonardo architecture.
Each MareNostrum 5 (and Leonardo) node has four GPUs, fully connected to one another. Our hierarchical Bine implementation exploits this topology by first performing an intra-node reduce-scatter, where each GPU exchanges data concurrently with the other three GPUs on the same node. Next, an inter-node Bine allreduce is executed among ranks with the same GPU local identifier, operating on the buffer segment received during the reduce-scatter phase. Finally, an intra-node allgather reconstructs the complete result, following a similar pattern to the initial reduce-scatter. As for NCCL, vectors are aggregated on the GPUs.
We evaluated Bine against both Open MPI and NCCL (v2.20.5 on MareNostrum 5 and v2.22.3 on Leonardo). On MareNostrum 5, NCCL outperforms MPI algorithms in most cases. However, our MPI implementation of Bine surpasses NCCL’s best-performing algorithm across all configurations from 16 to 256 GPUs, for vector sizes larger than 4 MiB. On average, Bine achieves a 5% improvement, with gains reaching up to 24% on 256 GPUs. On Leonardo, on 512 GPUs Bine improves performance by 15% compared to MPI and remains within 7% of NCCL.
7. Related Work
Locality-Aware and Hierarchical Algorithm Locality-aware, and hierarchical or multi-leader algorithms (Bienz et al., 2022; Subramoni et al., 2011; Bayatpour et al., 2017; Hidayetoglu et al., 2024) have been proposed for various collectives. These algorithms, however, either require knowledge of the underlying topology (Subramoni et al., 2011) or make assumptions about the partitioning of ranks across groups (Bienz et al., 2022). Unlike hierarchical methods, Bine trees are as generic as traditional collective algorithms and, as shown in Sec. 5.3.2, can also be effectively integrated into hierarchical algorithms.
Automatic Algorithm Synthesis Some methods aim to automatically determine the optimal collective algorithm based on network topology and collective specifications (Shah et al., 2022; Cai et al., 2021; Won et al., 2024). However, their complexity increases exponentially as the system scales. For example, computing a solution for 128 nodes can take up to 11 hours (Shah et al., 2022), and a new solution may need to be recalculated for each allocation, which makes these approaches not practical on shared supercomputers, where users do not have control over allocations.
Topology-Optimized Algorithms Researchers developed collective algorithms tailored to specific network topologies (Almási et al., 2005; Ma et al., 2021; Feng et al., 2022). In contrast, Bine trees are versatile and topology-agnostic and consistently enhance performance across a variety of topologies.
8. Conclusions
In this paper, we introduced Bine trees, a novel approach to constructing binomial trees that reduces the distance between communicating ranks and, as a result, reduces traffic on global links and improves performance. We designed new algorithms for eight different collective operations, outperforming state-of-the-art algorithms, including traditional binomial trees and topology-specific methods. We observed performance improvements of up to $5\times$ on $8192$ nodes and reductions in global link traffic of up to 33%, highlighting the potential of Bine to deliver significant gains in performance and efficiency across different network topologies.
The main strength of Bine trees lies in their generality, being as generic as standard binomial trees and not relying on specific assumptions about the underlying network topology or job allocation. This makes it well-suited for large-scale, shared supercomputing environments. In conclusion, thanks to their broad applicability and ability to reduce network traffic across global links, Bine trees represent a significant step toward more efficient, scalable, and flexible collective algorithms. They can effectively replace binomial trees and butterfly-based approaches, especially at large scales.
Acknowledgements.
We thank Daniele Di Bari (CINECA) for supporting us in executing large-scale tests on Leonardo, Kim McMahon (HPE Cray) for the support with Cray MPICH collective algorithms selection, Kurt Lust for the information on mapping of nodes to groups on LUMI, and Faveo Hörold (ETH Zürich and RIKEN) for helping with preliminary tests on Fugaku. This work is supported by the European Union Commission’s Horizon Europe, under grant agreements 101175702 (NET4EXA), by the Sapienza University Grants ADAGIO and D2QNeT (Bando per la ricerca di Ateneo 2023 and 2024), and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement PSAP, No. 101002047). Daniele De Sensi is a member of Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM). The authors used ChatGPT to assist with editing. All ideas, content, and conclusions are the authors’ own.